OpenStudy (anonymous):

Help with 6^(2x=7) = 12 ?

OpenStudy (dan815):

1 too many equal sign in there

OpenStudy (anonymous):

my bad 2x+7

OpenStudy (dumbcow):

logs :) take log of both sides then take exponent and put it out front thx to log property log(x^n) = n*log(x)

OpenStudy (anonymous):

\[6^{2x+7} = 12\]

OpenStudy (anonymous):

so log 6 and log 12 ?

OpenStudy (dumbcow):

yes

OpenStudy (isaiah.feynman):